# One-Shot Broadcast Joint Source-Channel Coding with Codebook Diversity

### v1

@@ Appendix D Proof of Theorem 3
$\stackrel{{\scriptstyle\text{(a)}}}{{\leq}}$ $\operatorname{\mathbb{E}}\bigl[\bigl(1+KP_{Z}(\mathcal{B}_{D}(W^{m}))2^{\iota_{X^{n};Y^{n}}(X^{n};Y^{n})}\bigr)^{-1}\bigr]$ (86)
$\leq$ $\operatorname{Pr}[-{\log}P_{Z}(\mathcal{B}_{D}(W^{m}))>\Omega]$
@@
$\{2^{-\Omega+\iota_{X^{n};Y^{n}}(X^{n};Y^{n})+\log K}\geq\sqrt{n}\}$ (89)
@@
$=$ $\varepsilon-\frac{\eta+B}{\sqrt{\min\{n,m\}}}$ (96)
@@
$P_{e}\leq\frac{\gamma}{\sqrt{m}}+\frac{1}{\sqrt{n}}+\varepsilon-\frac{\eta+B}{\sqrt{\min\{n,m\}}}.$ (97)
@@ Appendix E Proof of Theorem 4
$A_{i}=\begin{cases}\begin{bmatrix}\iota_{X;Y}(X_{i},Y_{1i})-C&\cdots&\iota_{X;Y}(X_{i};Y_{Li})-C\end{bmatrix}^{T},\\ \qquad 1\leq i\leq n\\ \begin{bmatrix}\jmath_{W}(W_{i},D)-R(D)&\cdots&\jmath_{W}(W_{i},D)-R(D)\end{bmatrix}^{T},\\ \qquad n+1\leq i\leq n+m.\end{cases}$ (102)

### v2

@@ Appendix D Proof of Theorem 3
$\stackrel{{\scriptstyle\text{(a)}}}{{\leq}}$ $\operatorname{\mathbb{E}}\bigl[\bigl(1+KP_{Z^{m}}(\mathcal{B}_{D}(W^{m}))2^{\iota_{X^{n};Y^{n}}(X^{n};Y^{n})}\bigr)^{-1}\bigr]$ (86)
$\leq$ $\operatorname{Pr}[-{\log}P_{Z^{m}}(\mathcal{B}_{D}(W^{m}))>\Omega]$
@@
$\{2^{\iota_{X^{n};Y^{n}}(X^{n};Y^{n})-\Omega+\log K}\geq\sqrt{n}\}$ (89)
@@
$=$ $\varepsilon+\frac{B-\eta}{\sqrt{\min\{n,m\}}}$ (96)
@@
$P_{e}\leq\frac{\gamma}{\sqrt{m}}+\frac{1}{\sqrt{n}}+\varepsilon+\frac{B-\eta}{\sqrt{\min\{n,m\}}}.$ (97)
@@ Appendix E Proof of Theorem 4
$A_{i}=\begin{cases}\begin{bmatrix}\iota_{X;Y}(X_{i},Y_{1i})-C&\cdots&\iota_{X;Y}(X_{i};Y_{Li})-C\end{bmatrix}^{T},\\ \qquad 1\leq i\leq n\\ \begin{bmatrix}R(D)-\jmath_{W}(W_{i-n},D)&\cdots&R(D)-\jmath_{W}(W_{i-n},D)\end{bmatrix}^{T},\\ \qquad n+1\leq i\leq n+m.\end{cases}$ (102)